Two blocks of mass m=25 kg and M=20 kg are connected via pulley with a
configuration as shown. Assuming that the surfaces are smooth and the angle θ=
45
. Find the acceleration blocks and tension of the cable.
1
Expert's answer
2020-10-05T10:53:20-0400


Apply Newton's second law for M:


"-Ma=T-Mg."

Apply Newton's second law for m:


"X: ma=T-mg\text{ sin}\theta."

The solution will be


"a=g\frac{M-m\text{ sin}\theta}{m+M}=0.67\text{ m/s}^2,\\\space\\
T=M(g-a)=182.6\text{ N}."